OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions

OpenStudy (anonymous):

need help

OpenStudy (ikram002p):

well try to solve it :D

OpenStudy (anonymous):

i did and i got it mixed some how with no solution and one solution

OpenStudy (ikram002p):

y = 3.5x - 3.5 y = -3.5x + 3.5 add up the equation then u would have 2y=0 y=0 then y = 3.5x - 3.5 0= 3.5x - 3.5 so =1 so there is a solution which is (1,0)

OpenStudy (ikram002p):

and it unique , so only one solution

OpenStudy (ikram002p):

got it ?

OpenStudy (anonymous):

ya i think so

OpenStudy (ikram002p):

:D ok